OpenStudy (anonymous):

for the derivative of y=ln(sin^2x) do I apply chain twice?

OpenStudy (anonymous):

does it help to know that \[\ln(\sin^2(x))=2\ln(\sin(x))\]?

OpenStudy (anonymous):

ohh... yeah, taking the exponent out is nice:) But just asking, would it be wrong (not only necessary, but also wrong) to apply chain twice?

OpenStudy (anonymous):

I mean not only NOT necessary...

OpenStudy (anonymous):

try it and see what happens you will get a longer expression, then some trig identity will probably give you the shorter version

OpenStudy (anonymous):

y=ln(( sinx)^2) y' = csc^2 x cosx

OpenStudy (anonymous):

no, hold... I am wrong

OpenStudy (anonymous):

sin^2x = (sinx)^2 d/dx = 2sinx cosx then ln((sinx)^2) = csc^2x * 2sinx cosx = cot x right?

OpenStudy (anonymous):

2 cotx (sorry)

OpenStudy (aum):

You got 2cot(x) either way.

OpenStudy (anonymous):

cool... just wierd to apply the chain twice... tnx for the help everyone!